Jiminy’s Cricket Farm issued a bond with 25 years to maturity and a semiannual coupon rate of 4 percent 3 years ago. The bond currently sells for 108 percent of its face value. The company’s tax rate is 22 percent.

What is the pretax cost of debt?

What is the aftertax cost of debt?
Answer #1

Years to maturity =25-3 =22 years

Pretax cost of debt is calculated using the RATE function as follows:

=RATE(nper,pmt,pv,fv)

=RATE(22*2,4%/2*100,-108,100)*2

=3.48%

Post tax cost of debt:

=3.48%*(1-0.22)

=2.71%